OpenStudy (anonymous):

y = 3x + 7 y = x - 9

OpenStudy (anonymous):

if y = x - 9, then x - 9 = 3x + 7 -2x = 16 x = -8 y = -8 -9 y = -17 x = -8, y = -17

OpenStudy (anonymous):

thank you love